# Evolving Collective Cognition in Human–Agent Hybrid Societies: How Agents Form Stances and Boundaries

### v1

@@ 5.1.5 Time and Cost Analysis
To assess the practicality of CMASE, human evaluators engaged in real-time interactions within environments containing varying numbers of agents and rated the system’s responsiveness on a 7-point Likert scale. Responses were generated using 5 different LLM servers. The results are shown in Figure 5.
@@
Figure 5: CMASE performance under different agent counts. (a) Duration per time step. (b) Human evaluator responsiveness ratings.
@@
Figure 5(a) illustrates that the time required for a single time step in CMASE grows linearly with the number of agents. With 1 to 10 agents, the average duration is 23s, which includes both the mean API call latency and the environment’s action execution time. It increases to 498s when 1,000 agents are present. Figure 5(b) presents boxplots of human evaluator scores. Human evaluators spend on average approximately 50s per time step (from receipt of perceptual information through action selection and goal input). As long as agent counts remain at or below 100, evaluators exhibit low sensitivity to system responsiveness. Once the duration of a single time step surpasses their 50s operation window, waiting time accrues after action submission and sensitivity to responsiveness is markedly heightened.
In terms of cost, each agent’s action in one time step incurs approximately \$0.10 in API expenses. By comparison, Park et al. [28] employed GPT-3.5-turbo to simulate two days of gameplay (approximately 192 to 576 time steps) for 25 generative agents, spending several thousand dollars and several days to complete their simulation.
5.2 Study 1
5.2.1 Goal and procedure
We selected studies published in top journals and replicated their survey results under comparable conditions. To ensure that the LLM had not already been trained on the studies being replicated, we employed the GPT-4o “2024-12-01-preview” model and limited our replication to papers published after that date.
Lee and Han [16] analyzed interview data collected in South Korea during the COVID-19 pandemic and demonstrated that regions with higher vegetation cover exhibited lower levels of social fragmentation, indicating stronger community bonds. Based on the descriptions in the original study, we used the environment maker module to construct maps of South Korean regions with varying degrees of vegetation cover. Because the original paper emphasized South Korea’s community isolation policies during the pandemic, we designed isolated community environments and placed 10 agents in each.
5.2.2 Measures
Once the simulation reached a quasi-steady state, we interviewed the agents to measure their sense of social fragmentation, which comprises three dimensions: distrust, exploitation and indifference. In addition, agents completed a 7-point Likert scale questionnaire corresponding to these three subdimensions. The subdimensions are defined as follows: (a) Distrust denotes skepticism or lack of confidence in the integrity and functionality of social structures. (b) Exploitation refers to the perception or experience of being used unjustly or immorally for another’s benefit in a fragmented society. (c) Indifference manifests as a lack of interest, concern, or empathy for social issues or the plight of others. After running the simulation for 50 time steps, we interviewed each agent and asked them to complete the questionnaire.
5.2.3 Analysis
@@
Such behavior reduced cognitive load from external stimuli and deepened her processing of task-relevant cues within a monotonous environment. These observations indicate that agents in creative professions (e.g., fashion designers) and those in analytical roles (e.g., lawyers) rely more heavily on orderly, low-stimulation interiors to anchor their cognition and preserve a stable, idiosyncratic mental world. Consequently, the stress-reduction benefit of green areas on their social fragmentation is comparatively muted.

### v2

@@ 4.6. Time and Cost Analysis
To assess the practicality of CMASE, we conducted a human-in-the-loop responsiveness evaluation. Five evaluators were recruited for this study. The selection criteria required participants to familiar with tabletop role-playing games to ensure they could accurately judge the system’s fluidity against genre standards. The evaluators engaged in real-time interactions within environments containing varying numbers of agents. They rated the system’s responsiveness on a 7-point Likert scale. The results are shown in Figure 3.
@@
Figure 3. CMASE performance under different agent counts. (a) Duration per time step. (b) Human evaluator responsiveness ratings.
@@
Figure 3(a) illustrates that the time required for a single time step in CMASE grows linearly with the number of agents. With 1 to 10 agents, the average duration is 23s, which includes both the mean API call latency and the environment’s action execution time. It increases to 498s when 1,000 agents are present.
Figure 3(b) presents boxplots of human evaluator scores. The purpose of this evaluation was not merely to assess speed, but to identify the critical threshold for real-time human-in-the-loop usability. Our findings reveal that human evaluators require an average of approximately 50s per time step to process information and select actions. Consequently, as long as the system’s calculation time remains below this 50s window (which corresponds to agent counts $\leq$ 100), the system latency is effectively masked by the human’s own cognitive load, resulting in stable, high responsiveness ratings.However, once the agent count exceeds 100, the system latency surpasses the human operation window, causing waiting time to accrue and responsiveness ratings to drop partially. At 500 agents, the responsiveness ratings hit the measurement floor. Testing beyond this threshold was deemed unnecessary due to the diminishing marginal utility of data in an already unacceptable latency range, indicating the practical upper bound for real-time human participation in CMASE.
In terms of cost, each agent’s action in one time step incurs approximately \$0.10 in API expenses. To put this into perspective, we compared our cost structure with the benchmark set by Park et al. (Park et al., 2023), who reported spending “several thousand dollars” to simulate 25 agents for two days (about 192–576 time steps). Applying CMASE’s rate to a simulation of the same scale would result in an estimated total cost of \$480 to \$1,440. This indicates that CMASE offers a more economically feasible solution for multi-agent simulations of this magnitude.
5. Conclusion
In this paper, we have introduced a novel comprehensive modeling framework called Computational Multi-Agents Society Experiments (CMASE) , which is featured in three key aspects, that is, real-time researcher embedding, highly customizable social environment construction, and an ethnography-inspired approach to virtual fieldwork. We conduct quantitative and qualitative experiments, including time and cost analysis alongside the replication of a field study on urban greenery and social fragmentation, to demonstrate the effectiveness and efficiency of our platform. We believe CMASE makes a significant step towards the next generation of human-AI collaborative sociological modeling. In the future, we plan to support larger-scale, multi-user real-time interventions, and will also incorporate more complex multimodal spatial information to further enhance the simulation of embodied human experiences and emergent social dynamics.
Appendix A Detailed Qualitative Analysis and Interview Transcripts
To investigate the behavioral mechanisms underlying the statistical reductions in social fragmentation, we adopted a virtual fieldwork approach, treating the simulation as an observable digital field. We conducted in-depth semi-structured interviews with agents to reconstruct their subjective experiences and decision-making rationales. Instead of random sampling, we employed a purposive sampling strategy, selecting representative cases from two contrasting behavioral clusters identified in the trajectory data: agents who consistently remained indoors and those who actively utilized the outdoor green spaces. The following analysis cites specific agents from these clusters to illustrate how their pre-defined personas interacted with the environment.
@@ Appendix A Detailed Qualitative Analysis and Interview Transcripts
Here, the term ”exploitation” refers explicitly to the specific Exploitation sub-dimension of social fragmentation defined in Section LABEL:subsec5.2.2 (the sense of being unjustly used or burdened), which manifests in this context as the agent’s internalized sense of professional struggle.
@@
Such behavior reduced cognitive load from external stimuli and deepened her processing of task-relevant cues within a monotonous environment. These observations suggest that agents assigned to simulate creative workflows (e.g., the fashion designer persona) and those emulating analytical roles (e.g., the lawyer persona) appear to prioritize orderly, low-stimulation interiors to anchor their simulated cognition and preserve a stable, idiosyncratic mental world. Consequently, within the scope of their defined personas, the stress-reduction benefit of green areas on their social fragmentation is comparatively muted.